# High Gain Fusion Target Design using Generative Artificial Intelligence

### v3

@@ II The Ubuntu Fusion Target
For brevity, we will describe the laser targets and drives. The Ubuntu Fusion Target is a simple, easy to fabricate, room temperature target. This is in contrast to the very impractical and complicated MagLIF DT ice burner, and current ICF target designs. As shown in Fig. 3, the target consists of a cylinder of beryllium, coated on the inside with solid fuel, and filled with DT gas. The target is driven by four lasers whose illumination patterns on the cylinder are shaped with phase plates, as shown in Fig. 4, either to form a Z-pinch or an RFP. The pattern of the laser drive effectively makes the target plasma into a substantial MIFED coil. The implosion is not driven by radiation pressure nor ablation pressure, but by magnetic pressure. Unlike pulsed-power, which must physically touch the plasma, the lasers stand off from the target and only require four small lines of sight in a plane perpendicular to the axial direction. Ubuntu genAI will be used to optimize the target and drive design, including the axial structure. It will also be used to ponderomotively stabilize the metastable stagnation.
@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [23]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [24]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\text{i}\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [25]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [26, 27]; the Mayer Cluster Expansion [28]; or the functional Taylor expansion of the action functional. This enables the unification of the four forces as a question of Lie group symmetries, that is a Grand Unified Theory (GUT) of fields, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [29] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the infinitesimal Lie group generator, for Einstein’s General Theory of Relativity [30], in 1958 [31].

### v4

@@ II The Ubuntu Fusion Target
For brevity, we will describe the laser targets and drives. The Ubuntu Fusion Target is a simple, easy to fabricate, room temperature target. This is in contrast to the very impractical and complicated MagLIF DT ice burner, and current ICF target designs. As shown in Fig. 3, the target consists of a cylinder of beryllium, coated on the inside with solid fuel, and filled with DT gas. The target is driven by four lasers whose illumination patterns on the cylinder are shaped with phase plates, as shown in Fig. 4, either to form a Z-pinch or an RFP. The pattern of the laser drive effectively makes the target plasma into a substantial MIFED coil. The implosion is not driven by radiation pressure nor ablation pressure, but by magnetic pressure which increases as the target is imploded. Unlike pulsed-power, which must physically touch the plasma, the lasers stand off from the target and only require four small lines of sight in a plane perpendicular to the axial direction. Ubuntu genAI will be used to optimize the target and drive design, including the axial structure. It will also be used to ponderomotively stabilize the metastable stagnation.
@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [23]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [24]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\text{i}\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [25]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [26, 27]; the Mayer Cluster Expansion [28]; or the functional Taylor expansion of the action functional, $\delta^{m}S/\delta f^{m}$. This enables the unification of the four forces as a question of Lie group symmetries, that is a Grand Unified Theory (GUT) of fields, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [29] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the infinitesimal Lie group generator, for Einstein’s General Theory of Relativity [30], in 1958 [31].